OpenStudy (anonymous):

x+ 2y = 10 x + y = 6 whats my answer? (0,-4) (2,2) (2,4) (-4,-2)

OpenStudy (phi):

can you subtract the two equations? try top equation minus bottom equation. what do you get ?

OpenStudy (anonymous):

(2,4)?????

OpenStudy (phi):

try top equation minus bottom equation. what do you get ? you get a new equation, not a number.

OpenStudy (anonymous):

OHH....

OpenStudy (phi):

like this start with x+ 2y = 10 x + y = 6 on the left side do x+2y -x - y (we subtract all the terms of the bottom equation) on the right side of the = sign we do 10 -6 (which is 4) now you have x+2y -x - y = 4 can you simplify the left side ?